# Well-posedness of McKean-Vlasov generalized multivalued BSDEs

### v1

@@ 2.1. Assumptions and results
    There exist $\lambda>3+16L^{2}$ and $\mu>2+4K^{2}$,
@@
    $(Y_{t},U_{t})\in\partial\varphi,$ $d\mathbb{P}\times dt$,  $(Y_{t},V_{t})\in\partial\psi,$ $d\mathbb{P}\times dA_{t}$  a.s. on $[0,T]$.
@@
$2\langle|Y_{s}-\tilde{Y}_{s}|,L\|Z_{s}-\tilde{Z}_{s}\|\rangle\leq 4L^{2}|Y_{s}-\tilde{Y}_{s}|^{2}+\frac{1}{4}\|Z_{s}-\tilde{Z}_{s}\|^{2}.$
@@ 2.2. Existence and uniqueness
Proposition 4: Let Assumptions I-V be satisfied. Then there exists a unique solution $(Y,Z,U,V)$ $(\mathcal{S}_{k}^{\lambda,\mu}\cap\mathcal{H}_{k}^{\lambda,\mu}\cap\widetilde{\mathcal{H}}_{k}^{\lambda,\mu})\times\mathcal{H}_{k\times d}^{\lambda,\mu}$ for (9).
@@ 2.2.1. A priori estimate
Given $\nabla\varphi_{\varepsilon}=\dfrac{Y_{s}^{\delta}-J_{\varepsilon}(Y_{s}^{\varepsilon})}{\varepsilon}$ and using the monotonicity of operator $\partial\varphi$ as discussed in [12], we obtain
@@ 2.2.2. Proof of Theorem 3
$\exists Y\in S_{k}^{\lambda,\mu}\cap H_{k}^{\lambda,\mu}\cap\tilde{H}_{k}^{\lambda,\mu},\quad\exists Z\in H_{k\times d}^{\lambda,\mu},$
$\lim_{\varepsilon\searrow 0}Y^{\varepsilon}=Y\quad\text{in }S_{k}^{\lambda,\mu}\cap H_{k}^{\lambda,\mu}\cap\tilde{H}_{k}^{\lambda,\mu},$
$\lim_{\varepsilon\searrow 0}Z^{\varepsilon}=Z\quad\text{in }H_{k\times d}^{\lambda,\mu}.$
@@
Hence there exist $U\in H_{k}^{\lambda,\mu}$ and $V\in\tilde{H}_{k}^{\lambda,\mu}$ such that for a subsequence $\varepsilon_{n}\searrow 0$,
$U^{\varepsilon_{n}}\rightharpoonup U,\quad\text{weakly in Hilbert space }H_{k}^{\lambda,\mu},$
$V^{\varepsilon_{n}}\rightharpoonup V,\quad\text{weakly in Hilbert space }\tilde{H}_{k}^{\lambda,\mu}.$
@@ 3. BSVIs with locally Lipschitz and non-linear growth conditions
    $\forall N\in\mathcal{N},\exists r_{N}\in\mathbb{R}$ such that, for any $y,y^{\prime},z,z^{\prime}$ satisfying $|y|,|y^{\prime}|,|z|,|z^{\prime}|\leq N$ and for any $\Pi,\Pi^{\prime}$ satisfying $W_{2}(\Pi,\delta_{0}),W_{2}(\Pi^{\prime},\delta_{0})\leq N$,
@@
    $|F_{n}(t,y,z,\Pi)-F_{n}(t,y^{\prime},z^{\prime},\Pi^{\prime})|\leq L(|y-y^{\prime}|+|z-z^{\prime}|+W_{2}(\Pi,\Pi^{\prime})),$
@@
    $\rho_{N}^{2}(f):=\mathbb{E}\left[\int_{0}^{T}\sup_{|y|,|z|,W_{2}(\Pi,\delta_{0})\leq N}|f(t,y,z,\Pi)|^{2}\,\mathrm{d}t\right].$
@@
Analogous to the proof of Proposition 4, we have
@@
$\mathbb{E}\int_{0}^{T}e^{\lambda s+\mu A_{s}}\left(|\nabla\varphi_{\frac{1}{n}}(Y_{s}^{\frac{1}{n}})|^{2}\mathrm{d}s+|\nabla\psi_{\frac{1}{n}}(Y_{s}^{\frac{1}{n}})|^{2}\mathrm{d}A_{s}\right)\leq C\mathcal{M}(T),$
$\mathbb{E}\int_{0}^{T}e^{\lambda s+\mu A_{s}}\left(\varphi(J_{\frac{1}{n}}(Y_{s}^{\frac{1}{n}}))\mathrm{d}s+\psi(\hat{J}_{\frac{1}{n}}(Y_{s}^{\frac{1}{n}}))\mathrm{d}A_{s}\right)\leq C\mathcal{M}(T),$
$\mathbb{E}e^{\lambda t+\mu A_{t}}\left(|Y_{t}^{\frac{1}{n}}-J_{\frac{1}{n}}(Y_{t}^{\frac{1}{n}})|^{2}+|Y_{t}^{\frac{1}{n}}-\hat{J}_{\frac{1}{n}}(Y_{t}^{\frac{1}{n}})|^{2}\right)\leq{\frac{C}{n}}\mathcal{M}(T),$
$\mathbb{E}e^{\lambda t+\mu A_{t}}\left(\varphi(J_{\frac{1}{n}}(Y_{t}^{\frac{1}{n}}))+\psi(\hat{J}_{\frac{1}{n}}(Y_{t}^{\frac{1}{n}}))\right)\leq C\mathcal{M}(T).$
Step 3): For any given $N$, we set $D_{n,m}^{N}:=\{(s,w):|Y_{s}^{\frac{1}{n}}|^{2}+|Y_{s}^{\frac{1}{m}}|^{2}+|Z_{s}^{\frac{1}{n}}|^{2}+|Z_{s}^{\frac{1}{m}}|^{2}\geq N^{2}\}$ and denote $\bar{D}_{n,m}^{N}=\Omega-D_{n,m}^{N}$.
$\mathbb{E}e^{\lambda t+\mu A_{t}}|Y_{t}^{\frac{1}{n}}-Y_{t}^{\frac{1}{m}}|^{2}+\mathbb{E}\int_{t}^{T}e^{\lambda s+\mu A_{s}}|Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}}|^{2}(\lambda\mathrm{d}s+\mu\mathrm{d}A_{s})$
$\quad+\mathbb{E}\int_{t}^{T}e^{\lambda s+\mu A_{s}}\|Z_{s}^{\frac{1}{n}}-Z_{s}^{\frac{1}{m}}\|^{2}\mathrm{d}s$
$\quad+2\mathbb{E}\int_{t}^{T}e^{\lambda s+\mu A_{s}}\langle Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}},\nabla\varphi_{n}(Y_{s}^{\frac{1}{n}})-\nabla\varphi_{m}(Y_{s}^{\frac{1}{m}})\rangle\mathrm{d}s$
$\quad+2\mathbb{E}\int_{t}^{T}e^{\lambda s+\mu A_{s}}\langle Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}},\nabla\psi_{n}(Y_{s}^{\frac{1}{n}})-\nabla\psi_{m}(Y_{s}^{\frac{1}{m}})\rangle\mathrm{d}A_{s}$
$=2\mathbb{E}\int_{t}^{T}e^{\lambda s+\mu A_{s}}\langle Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}},F_{n}(s,Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}},\mathcal{L}(Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}}))-F_{m}(s,Y_{s}^{\frac{1}{m}},Z_{s}^{\frac{1}{m}},\mathcal{L}(Y_{s}^{\frac{1}{m}},Z_{s}^{\frac{1}{m}}))\rangle\mathrm{d}s$
$\quad+2\mathbb{E}\int_{t}^{T}e^{\lambda s+\mu A_{s}}\langle Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}},G(s,Y_{s}^{\frac{1}{n}})-G(s,Y_{s}^{\frac{1}{m}})\rangle\mathrm{d}A_{s}$
$=\Gamma_{1}+\Gamma_{2}+\Gamma_{3}+\Gamma_{4}++\Gamma_{5},$
@@
$\Gamma_{1}=2\mathbb{E}\int_{t}^{T}e^{\lambda s+\mu A_{s}}\langle Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}},F_{n}(s,Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}},\mathcal{L}(Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}}))-F_{m}(s,Y_{s}^{\frac{1}{m}},Z_{s}^{\frac{1}{m}},\mathcal{L}(Y_{s}^{\frac{1}{m}},Z_{s}^{\frac{1}{m}}))\rangle 1_{D_{n,m}^{N}}\mathrm{d}s,$
$\Gamma_{2}=2\mathbb{E}\int_{t}^{T}e^{\lambda s+\mu A_{s}}\langle Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}},F_{n}(s,Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}},\mathcal{L}(Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}}))-F(s,Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}},\mathcal{L}(Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}}))\rangle 1_{\bar{D}_{n,m}^{N}}\mathrm{d}s,$
$\Gamma_{3}=2\mathbb{E}\int_{t}^{T}e^{\lambda s+\mu A_{s}}\langle Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}},F(s,Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}},\mathcal{L}(Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}}))-F(s,Y_{s}^{\frac{1}{m}},Z_{s}^{\frac{1}{m}},\mathcal{L}(Y_{s}^{\frac{1}{m}},Z_{s}^{\frac{1}{m}}))\rangle 1_{\bar{D}_{n,m}^{N}}\mathrm{d}s,$
$\Gamma_{4}=2\mathbb{E}\int_{t}^{T}e^{\lambda s+\mu A_{s}}\langle Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}},F(s,Y_{s}^{\frac{1}{m}},Z_{s}^{\frac{1}{m}},\mathcal{L}(Y_{s}^{\frac{1}{m}},Z_{s}^{\frac{1}{m}}))-F_{m}(s,Y_{s}^{\frac{1}{m}},Z_{s}^{\frac{1}{m}},\mathcal{L}(Y_{s}^{\frac{1}{m}},Z_{s}^{\frac{1}{m}}))\rangle 1_{\bar{D}_{n,m}^{N}}\mathrm{d}s,$
$\Gamma 5=2\mathbb{E}\int_{t}^{T}e^{\lambda s+\mu A_{s}}\langle Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}},G(s,Y_{s}^{\frac{1}{n}})-G(s,Y_{s}^{\frac{1}{m}})\rangle\mathrm{d}A_{s}.$
Utilizing the assumptions, properties of $D_{n,m}^{N}$ and (18), we derive the following inequalities.
$\Gamma_{2}+\Gamma_{4}\leq 2\mathbb{E}\int_{t}^{T}e^{\lambda s+\mu A_{s}}|Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}}|^{2}\mathrm{d}s+\rho_{N}(F_{n}-F)+\rho_{N}(F_{m}-F),$
$\Gamma_{5}\leq(K^{2}+1)\mathbb{E}\int_{t}^{T}e^{\lambda s+\mu A_{s}}|Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}}|^{2}\mathrm{d}A_{s},$
[…]
$\quad+\mathbb{E}\int_{t}^{T}e^{\lambda s+\mu A_{s}}\big[\mathbb{E}|Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}}|^{2}+\frac{1}{4}\mathbb{E}\|Z_{s}^{\frac{1}{n}}-Z_{s}^{\frac{1}{m}}\|^{2}\big]1_{\bar{D}_{n,m}^{N}}\mathrm{d}s$
@@
$\qquad\times\dfrac{|Y_{s}^{\frac{1}{n}}|^{2}+|Y_{s}^{\frac{1}{m}}|^{2}+|Z_{s}^{\frac{1}{n}}|^{2}+|Z_{s}^{\frac{1}{m}}|^{2}}{N^{2}}1_{D_{n,m}^{N}}\mathrm{d}s$
@@
$\qquad\times\dfrac{\big(|Y_{s}^{\frac{1}{n}}|^{2}+|Y_{s}^{\frac{1}{m}}|^{2}+|Z_{s}^{\frac{1}{n}}|^{2}+|Z_{s}^{\frac{1}{m}}|^{2}\big)^{1-\alpha}}{N^{2(1-\alpha)}}1_{D_{n,m}^{N}}\mathrm{d}s$
$\leq\mathbb{E}\int_{t}^{T}e^{\lambda s+\mu A_{s}}|Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}}|^{2}\mathrm{d}s+\frac{C}{N^{2(1-\alpha)}}\mathcal{M}(T)$
$\quad+\frac{20L^{2}}{N^{2}}\mathbb{E}\int_{t}^{T}\mathbb{E}\big[\sup_{0\leq s\leq T}e^{\lambda s+\mu A_{s}}(|Y_{s}^{\frac{1}{n}}|^{2}+|Y_{s}^{\frac{1}{m}}|^{2})+\frac{1}{T}\int_{0}^{T}e^{\lambda s+\mu A_{s}}(\|Z_{s}^{\frac{1}{n}}\|^{2}+\|Z_{s}^{\frac{1}{m}}\|^{2}+\eta_{s}^{2})\big]$
$\qquad\times e^{\lambda s+\mu A_{s}}(|Y_{s}^{\frac{1}{n}}|^{2}+|Y_{s}^{\frac{1}{m}}|^{2}+|Z_{s}^{\frac{1}{n}}|^{2}+|Z_{s}^{\frac{1}{m}}|^{2})\mathrm{d}s$
$\leq\mathbb{E}\int_{t}^{T}e^{\lambda s+\mu A_{s}}|Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}}|^{2}\mathrm{d}s+\frac{C}{N^{2(1-\alpha)}}\mathcal{M}(T)+\frac{20L^{2}}{N^{2}}(3+\frac{2}{T})\mathcal{M}^{2}(T)$
$\leq\mathbb{E}\int_{t}^{T}e^{\lambda s+\mu A_{s}}|Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}}|^{2}\mathrm{d}s+\frac{C}{N^{2(1-\alpha)}}[\mathcal{M}(T)+\mathcal{M}^{2}(T)].$
Summarizing the above and taking into account the ranges of $\lambda$ and $\mu$, we have
$\mathbb{E}e^{\lambda t+\mu A_{t}}|Y_{t}^{\frac{1}{n}}-Y_{t}^{\frac{1}{m}}|^{2}+\frac{1}{2}\mathbb{E}\int_{t}^{T}e^{\lambda s+\mu A_{s}}\|Z_{s}^{\frac{1}{n}}-Z_{s}^{\frac{1}{m}}\|^{2}\mathrm{d}s$

### v2

@@ 2.1. Assumptions and results
    There exist $\lambda>4+16L^{2}$ and $\mu>2+4K^{2}$,
@@
    $(Y_{t},U_{t})\in\partial\varphi,$ $d\mathbb{P}\times dt$,  $(Y_{t},V_{t})\in\partial\psi,$ $d\mathbb{P}\times dA_{t}$  a.s. on $[0,T]$, i.e
    $\langle U_{t},u_{t}-Y_{t}\rangle+\varphi(Y_{t})\leq\varphi(u_{t}),\,\quad\forall u\in\mathcal{H}_{k}^{\lambda,\mu},\quad\langle V_{t},v_{t}-Y_{t}\rangle+\varphi(Y_{t})\leq\varphi(v_{t}),\,\quad\forall v\in\tilde{\mathcal{H}}_{k}^{\lambda,\mu},$
@@
$2L|Y_{s}-\tilde{Y}_{s}|\|Z_{s}-\tilde{Z}_{s}\|\leq 4L^{2}|Y_{s}-\tilde{Y}_{s}|^{2}+\frac{1}{4}\|Z_{s}-\tilde{Z}_{s}\|^{2}.$
@@
Proposition 4: Let Assumptions I-V be satisfied. Then there exists a unique solution $(Y,Z,U,V)\in(\mathcal{S}_{k}^{\lambda,\mu}\cap\mathcal{H}_{k}^{\lambda,\mu}\cap\widetilde{\mathcal{H}}_{k}^{\lambda,\mu})\times\mathcal{H}_{k\times d}^{\lambda,\mu}\times\mathcal{H}_{k}^{\lambda,\mu}\times\tilde{\mathcal{H}}_{k}^{\lambda,\mu}$ for (9).
@@ 2.2.1. A priori estimate
Given $\nabla\varphi_{\varepsilon}=\dfrac{Y_{s}^{\varepsilon}-J_{\varepsilon}(Y_{s}^{\varepsilon})}{\varepsilon}$ and using the property of operator $\partial\varphi$ as discussed in [12], which is $\nabla\varphi_{\varepsilon}(Y_{t}^{\varepsilon})\in\partial\varphi(J_{\varepsilon}(Y_{t}^{\varepsilon}))$, we obtain
@@ 2.2.2. Proof of Theorem 3
$\exists Y\in\mathcal{S}_{k}^{\lambda,\mu}\cap H_{k}^{\lambda,\mu}\cap\tilde{\mathcal{H}}_{k}^{\lambda,\mu},\quad\exists Z\in\mathcal{H}_{k\times d}^{\lambda,\mu},$
$\lim_{\varepsilon\searrow 0}Y^{\varepsilon}=Y\quad\text{in }\mathcal{S}_{k}^{\lambda,\mu}\cap\mathcal{H}_{k}^{\lambda,\mu}\cap\tilde{\mathcal{H}}_{k}^{\lambda,\mu},$
$\lim_{\varepsilon\searrow 0}Z^{\varepsilon}=Z\quad\text{in }\mathcal{H}_{k\times d}^{\lambda,\mu}.$
@@
Hence there exist $U\in\mathcal{H}_{k}^{\lambda,\mu}$ and $V\in\tilde{\mathcal{H}}_{k}^{\lambda,\mu}$ such that for a subsequence $\varepsilon_{n}\searrow 0$,
$U^{\varepsilon_{n}}\rightharpoonup U,\quad\text{weakly in Hilbert space }\mathcal{H}_{k}^{\lambda,\mu},$
$V^{\varepsilon_{n}}\rightharpoonup V,\quad\text{weakly in Hilbert space }\tilde{\mathcal{H}}_{k}^{\lambda,\mu}.$
@@ 3. BSVIs with locally Lipschitz and non-linear growth conditions
    $\forall N\in\mathcal{N},\exists r_{N}\in\mathbb{R}$ such that, for any $y,y^{\prime},z,z^{\prime}$ satisfying $|y|,|y^{\prime}|,\|z\|,\|z^{\prime}\|\leq N$ and for any $\Pi,\Pi^{\prime}$ satisfying $W_{2}(\Pi,\delta_{0}),W_{2}(\Pi^{\prime},\delta_{0})\leq N$,
@@
    $|F_{n}(t,y,z,\Pi)-F_{n}(t,y^{\prime},z^{\prime},\Pi^{\prime})|\leq L_{n}(|y-y^{\prime}|+|z-z^{\prime}|+W_{2}(\Pi,\Pi^{\prime})),$
@@
    $\rho_{N}^{2}(f):=\mathbb{E}\left[\int_{0}^{T}\sup_{|y|,|z|,W_{2}(\Pi,\delta_{0})\leq N}e^{\lambda t+\mu A_{t}}|f(t,y,z,\Pi)|^{2}\,\mathrm{d}t\right].$
@@
First, we present the proof of existence. Analogous to the proof of Proposition 4, we have
$2\langle Y_{s}^{\frac{1}{n}},G(s,Y_{s}^{\frac{1}{n}})\rangle\leq(2+K^{2})|Y_{s}^{\frac{1}{n}}|^{2}+\nu_{s}^{2}.\qquad\qquad\qquad\qquad\qquad$
Considering the estimates of the generator $F$ in its integral form, we have
$2\int_{t}^{T}e^{\lambda s+\mu A_{s}}\langle Y_{s}^{\frac{1}{n}},F_{n}(s,Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}},\mathcal{L}(Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}}))\rangle\mathrm{d}s$
$\leq\int_{t}^{T}e^{\lambda s+\mu A_{s}}\big[p|Y_{s}^{\frac{1}{n}}|^{2}+\frac{1}{p}|F_{n}(s,Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}},\mathcal{L}(Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}}))|^{2}\big]\mathrm{d}s$
$\leq\int_{t}^{T}e^{\lambda s+\mu A_{s}}\big[p|Y_{s}^{\frac{1}{n}}|^{2}+\frac{1}{p}|L(|Y_{s}^{\frac{1}{n}}|^{\alpha}+\|Z_{s}^{\frac{1}{n}}\|^{\alpha}+W_{2}(\Pi,\delta_{0}))+\eta_{t}|^{2}\big]\mathrm{d}s$
$\leq\int_{t}^{T}e^{\lambda s+\mu A_{s}}\big[p|Y_{s}^{\frac{1}{n}}|^{2}+\frac{1}{p}\big(qL^{2}(|Y_{s}^{\frac{1}{n}}|^{\alpha}+\|Z_{s}^{\frac{1}{n}}\|^{\alpha}+W_{2}(\Pi,\delta_{0}))^{2}+\frac{1}{q}\eta_{t}^{2}\big)\big]\mathrm{d}s$
$\leq\int_{t}^{T}e^{\lambda s+\mu A_{s}}\big[p|Y_{s}^{\frac{1}{n}}|^{2}+\frac{1}{pq}\eta_{t}^{2}+\frac{3L^{2}q}{p}(|Y_{s}^{\frac{1}{n}}|^{2\alpha}+\|Z_{s}^{\frac{1}{n}}\|^{2\alpha})+\frac{6L^{2}q}{p}(\mathbb{E}|Y_{s}^{\frac{1}{n}}|^{2}+\mathbb{E}\|Z_{s}^{\frac{1}{n}}\|^{2})\big]\mathrm{d}s.$
Applying the integral form of Hölder’s inequality to the terms involving $|Y_{s}^{\frac{1}{n}}|^{2\alpha}$, we obtain
$\int_{t}^{T}e^{\lambda s+\mu A_{s}}|Y_{s}^{\frac{1}{n}}|^{2\alpha}\mathrm{d}s$
$\leq\big[\int_{t}^{T}\big(e^{\alpha(\lambda s+\mu A_{s})}|Y_{s}^{\frac{1}{n}}|^{2\alpha}\big)^{\frac{1}{\alpha}}\mathrm{d}s\big]^{\alpha}\big[\int_{t}^{T}\big(e^{(1-\alpha)(\lambda s+\mu A_{s})}\big)^{\frac{1}{1-\alpha}}\mathrm{d}s\big]^{1-\alpha}$
$\leq\big[\int_{t}^{T}e^{\lambda s+\mu A_{s}}|Y_{s}^{\frac{1}{n}}|^{2}\mathrm{d}s\big]^{\alpha}\big(\int_{t}^{T}e^{\lambda s+\mu A_{s}}\mathrm{d}s\big)^{1-\alpha}$
$\leq Te^{\lambda T+\mu A_{T}}\int_{t}^{T}e^{\lambda s+\mu A_{s}}|Y_{s}^{\frac{1}{n}}|^{2}\mathrm{d}s.$
Following a similar procedure for the terms containing $|Z_{s}^{\frac{1}{n}}|^{2\alpha}$ and setting $q=\frac{1}{Te^{\lambda T+\mu A_{T}}},p=12L^{2}$, the previous expressions can be consolidated as follows:
$2\int_{t}^{T}e^{\lambda s+\mu A_{s}}\langle Y_{s}^{\frac{1}{n}},F_{n}(s,Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}},\mathcal{L}(Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}}))\rangle\mathrm{d}s$
@@
$\mathbb{E}\int_{0}^{T}e^{\lambda s+\mu A_{s}}\left(|\nabla\varphi_{\frac{1}{n}}(Y_{s}^{\frac{1}{n}})|^{2}\mathrm{d}s+|\nabla\psi_{\frac{1}{n}}(Y_{s}^{\frac{1}{n}})|^{2}\mathrm{d}A_{s}\right)\leq C_{n}\mathcal{M}(T),$
$\mathbb{E}\int_{0}^{T}e^{\lambda s+\mu A_{s}}\left(\varphi(J_{\frac{1}{n}}(Y_{s}^{\frac{1}{n}}))\mathrm{d}s+\psi(\hat{J}_{\frac{1}{n}}(Y_{s}^{\frac{1}{n}}))\mathrm{d}A_{s}\right)\leq C_{n}\mathcal{M}(T),$
$\mathbb{E}e^{\lambda t+\mu A_{t}}\left(|Y_{t}^{\frac{1}{n}}-J_{\frac{1}{n}}(Y_{t}^{\frac{1}{n}})|^{2}+|Y_{t}^{\frac{1}{n}}-\hat{J}_{\frac{1}{n}}(Y_{t}^{\frac{1}{n}})|^{2}\right)\leq{\frac{C_{n}}{n}}\mathcal{M}(T),$
$\mathbb{E}e^{\lambda t+\mu A_{t}}\left(\varphi(J_{\frac{1}{n}}(Y_{t}^{\frac{1}{n}}))+\psi(\hat{J}_{\frac{1}{n}}(Y_{t}^{\frac{1}{n}}))\right)\leq C_{n}\mathcal{M}(T).$
Step 3): For any given NN, we set $D_{n,m}^{N}:=\{(s,w):|Y_{s}^{\frac{1}{n}}|^{2}+|Y_{s}^{\frac{1}{m}}|^{2}+|Z_{s}^{\frac{1}{n}}|^{2}+|Z_{s}^{\frac{1}{m}}|^{2}\geq N^{2}\}$ and denote $\bar{D}_{n,m}^{N}=\Omega-D_{n,m}^{N}$.
By using Itô formula and taking intergal, we have
$e^{\lambda t+\mu A_{t}}|Y_{t}^{\frac{1}{n}}-Y_{t}^{\frac{1}{m}}|^{2}+\int_{t}^{T}e^{\lambda s+\mu A_{s}}|Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}}|^{2}(\lambda\mathrm{d}s+\mu\mathrm{d}A_{s})+\int_{t}^{T}e^{\lambda s+\mu A_{s}}\|Z_{s}^{\frac{1}{n}}-Z_{s}^{\frac{1}{m}}\|^{2}\mathrm{d}s$
$\quad+2\int_{t}^{T}e^{\lambda s+\mu A_{s}}\langle Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}},\nabla\varphi_{n}(Y_{s}^{\frac{1}{n}})-\nabla\varphi_{m}(Y_{s}^{\frac{1}{m}})\rangle\mathrm{d}s$
$\quad+2\int_{t}^{T}e^{\lambda s+\mu A_{s}}\langle Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}},\nabla\psi_{n}(Y_{s}^{\frac{1}{n}})-\nabla\psi_{m}(Y_{s}^{\frac{1}{m}})\rangle\mathrm{d}A_{s}$
$=2\int_{t}^{T}e^{\lambda s+\mu A_{s}}\langle Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}},F_{n}(s,Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}},\mathcal{L}(Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}}))-F_{m}(s,Y_{s}^{\frac{1}{m}},Z_{s}^{\frac{1}{m}},\mathcal{L}(Y_{s}^{\frac{1}{m}},Z_{s}^{\frac{1}{m}}))\rangle\mathrm{d}s$
$\quad+2\int_{t}^{T}e^{\lambda s+\mu A_{s}}\langle Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}},G(s,Y_{s}^{\frac{1}{n}})-G(s,Y_{s}^{\frac{1}{m}})\rangle\mathrm{d}A_{s}$
$\quad-2\int_{t}^{T}e^{\lambda s+\mu A_{s}}\langle\langle Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}},(Z_{s}^{\frac{1}{n}}-Z_{s}^{\frac{1}{m}})dW_{s}\rangle$
$=:\Gamma_{1}+\Gamma_{2}+\Gamma_{3}+\Gamma_{4}+\Gamma_{5}+\Gamma_{6},$
@@
$\Gamma_{1}=2\int_{t}^{T}e^{\lambda s+\mu A_{s}}\langle Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}},\Theta_{1}\rangle 1_{D_{n,m}^{N}}\mathrm{d}s,\quad\Gamma_{2}=2\int_{t}^{T}e^{\lambda s+\mu A_{s}}\langle Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}},\Theta_{2}\rangle 1_{\bar{D}_{n,m}^{N}}\mathrm{d}s,$
$\Gamma_{3}=2\int_{t}^{T}e^{\lambda s+\mu A_{s}}\langle Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}},\Theta_{3}\rangle 1_{\bar{D}_{n,m}^{N}}\mathrm{d}s,\quad\Gamma_{4}=2\int_{t}^{T}e^{\lambda s+\mu A_{s}}\langle Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}},\Theta_{4}\rangle 1_{\bar{D}_{n,m}^{N}}\mathrm{d}s,$
$\Gamma 5=2\int_{t}^{T}e^{\lambda s+\mu A_{s}}\langle Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}},G(s,Y_{s}^{\frac{1}{n}})-G(s,Y_{s}^{\frac{1}{m}})\rangle\mathrm{d}A_{s},$
$\Gamma_{6}=-2\int_{t}^{T}e^{\lambda s+\mu A_{s}}\langle\langle Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}},(Z_{s}^{\frac{1}{n}}-Z_{s}^{\frac{1}{m}})dW_{s}\rangle,$
$\Theta_{1}=F_{n}(s,Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}},\mathcal{L}(Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}}))-F_{m}(s,Y_{s}^{\frac{1}{m}},Z_{s}^{\frac{1}{m}},\mathcal{L}(Y_{s}^{\frac{1}{m}},Z_{s}^{\frac{1}{m}})),$
$\Theta_{2}=F_{n}(s,Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}},\mathcal{L}(Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}}))-F(s,Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}},\mathcal{L}(Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}})),$
$\Theta_{3}=F(s,Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}},\mathcal{L}(Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}}))-F(s,Y_{s}^{\frac{1}{m}},Z_{s}^{\frac{1}{m}},\mathcal{L}(Y_{s}^{\frac{1}{m}},Z_{s}^{\frac{1}{m}})),$
$\Theta_{4}=F(s,Y_{s}^{\frac{1}{m}},Z_{s}^{\frac{1}{m}},\mathcal{L}(Y_{s}^{\frac{1}{m}},Z_{s}^{\frac{1}{m}}))-F_{m}(s,Y_{s}^{\frac{1}{m}},Z_{s}^{\frac{1}{m}},\mathcal{L}(Y_{s}^{\frac{1}{m}},Z_{s}^{\frac{1}{m}})).$
Utilizing the assumptions, properties of $D_{n,m}^{N}$, (18) and ${x}^{\alpha}\leq 1+|x|,\alpha\in[0,1)$, we derive the following inequalities
@@
$\qquad+\int_{0}^{T}e^{\lambda r+\mu A_{r}}\|Z_{s}^{\frac{1}{m}}\|^{2}\mathrm{d}r\big)\dfrac{|Y_{s}^{\frac{1}{n}}|^{2}+|Y_{s}^{\frac{1}{m}}|^{2}+|Z_{s}^{\frac{1}{n}}|^{2}+|Z_{s}^{\frac{1}{m}}|^{2}}{N^{2}}1_{D_{n,m}^{N}}\mathrm{d}s$
@@
$\leq\int_{t}^{T}e^{\lambda s+\mu A_{s}}|Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}}|^{2}\mathrm{d}s$